whats the radius of a circle with the area of 192 cm is
Answer by stanbon(75887)

whats the radius of a circle with the area of 192 cm is
-----------------
Area = pi * r^2
192 = pi *r^2
r^2 = 192/pi = 61.12
r = 7.82 cm\
================
Cheers,
Stan H.